

********************************************************************
UARS WEEKLY STATUS REPORT - 22 Jun 2001
********************************************************************
***********
SOLSTICE
***********
Other than being turned off for a few hours yesterday (01172) due to the
solar eclipse low power situation, SOLSTICE continues to operate normally,
taking solar data on all available orbital tracking periods and stellar
calibration observations whenever possible. We continue to gather enough
data to meet all the SOLSTICE science objectives during this period of
reduced tape recorder operations.
***********
SUSIM
***********
On Thursday, June 21, SUSIM was switched off when a UARS TMON tripped.
The instrument was switched back on before the end of the day.
Otherwise, SUSIM completed its daily solar scans during this past week
during periods of predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI continues to function normally, making wind measurements during
the sunlit parts of orbits nearly every day, and measuring nighttime
winds when power is available.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 09 June 2001 (Orbit 53266, GMT Day 160) through 15 June 2001
(Orbit 53370, GMT Day 166).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 166/1315 53364 Power On
The beta angle ranged between the angles of 30.8 degrees to 33.4 degrees for
this report period. The beta angle is now decreasing towards a minimum angle
of 2 degrees on 28 June 2001 (DOY 179).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.6 to 1.9 degrees C. Additional battery performance data for each
battery on 8 June 2001 and 15 June 2001 are provided in Appendix D.
The clock error ranged from Ð12.0 to 12.6 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 24
¥ WINDII = 03
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 03
Four (4) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/119 - 2287.499116
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/158 show the observatory data loss to be 301 hours, 02
minutes, 12.728 seconds. This shows zero loss since the last report period.
This is a 0.39 percent data loss, which equals a 99.61 percent data capture for
the mission.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to fine-tune the TAM alignment by loading new alignment
matrices. A new alignment is required for the reverse flight direction. This
is a slow process and it will take many alignments to lock in on the most
accurate solution.
An instrument anomaly was experienced on 05 May 2001. The instrument involved
was the ACRIM II instrument. All ACRIM telemetry for currents and voltages all
dropped to zero values suddenly. The ACRIM team in coordination with the
UARSFOT is conducting analysis and command tests to determine the cause of the
anomaly. Preliminary analysis shows that the anomaly was caused due to failed
circuitry in the ACRIM electronics box. The FOT is awaiting a future plan of
action from the ACRIM team.
MEETINGS
The FOT attended a UARS FS CCB Meeting on Monday, June 11, 2001.
The FOT attended a UARS TAC RS Meeting on Wednesday, June 13, 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 28 June 2001 (DOY 179). This maneuver
will be a Reverse to Forward Flight maneuver (Southern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 15 June 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
53268 160/0357 01-377 OBC Clk Adj 20.6 to 0.0 msec/day
53269 160/0615 01-376 NBTR NOOPS D-160
53277 160/1819 01-378 NBTR NOOPS D-161
53281 161/0041 01-379 OBC Clk Adj 0.0 to -20.6 msec/day
53284 161/0538 01-378 NBTR NOOPS D-161
53292 161/1919 01-380 NBTR NOOPS D-162
53293 161/2100 01-381 OBC Clk Adj -20.6 to 0.0 msec/day
53297 162/0216 01-382 OBC Clk Adj 0.0 to 20.6 msec/day
53299 162/0539 01-380 NBTR NOOPS D-162
53303 162/1306 01-383 TAM/OBC TAM Alignment
53307 162/1821 01-384 NBTR NOOPS D-163
53314 163/0539 01-384 NBTR NOOPS D-163
53319 163/1041 01-385 OBC/CLK Clk Adj 20.6 to 0.0 msec/day
53329 164/0540 01-386 NBTR NOOPS D-164
53332 164/1122 01-386 NBTR PLAYBACK / RTS-58
53336 164/1740 01-386 NBTR NOOPS D-164
53337 164/1822 01-387 NBTR NOOPS D-165
53339 164/2101 01-388 OBC TABLE 65 (STAR CATALOG)
53344 165/0540 01-387 NBTR NOOPS D-165
53351 165/1641 01-387 NBTR NOOPS D-165
53352 165/1823 01-389 NBTR NOOPS D-166
53358 166/0439 01-389 NBTR NOOPS D166
53364 166/1315 01-390 WINDII Power On
53365 166/1507 01-391 WINDII DAY 166 MICROLOAD AS REQUESTED
53366 166/1741 01-389 NBTR NOOPS D-166
53367 166/1823 01-393 NBTR NOOPS D-167
53367 166/1921 01-394 WINDII Reload instrument
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 15 June 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 15 June 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
53268 160/0401 Clk Adj 20.6 to 0.0 msec/day
53281 161/0044 Clk Adj 0.0 to -20.6 msec/day
53293 161/2104 Clk Adj -20.6 to 0.0 msec/day
53297 162/0232 Clk Adj 0.0 to 20.6 msec/day
53319 163/1050 Clk Adj 20.6 to 0.0 msec/day
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 15 June 2001
8 June 2001 - GMT Day 159 Beta = 26.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 12.0 0.000 0.0
2 +11.2/-112.0 +6.44/+4.51 +28.8/-15.2 23.8 1.030 21.9
3 +22.4/-67.2 +5.47/+3.23 +28.4/-16.0 23.8 1.020 21.6
15 June 2001 - GMT Day 166 Beta = 33.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 12.0 0.000 0.0
2 +11.2/-78.4 +6.44/+4.51 +28.4/-15.2 23.8 1.030 20.2
3 +16.8/-89.6 +5.15/+3.55 +28.0/-15.2 23.8 1.020 19.9
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
---- ---- ---------
-no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 08 June 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-032 FOT 53272 160/1121 14170 Generic Late Acquisition
01-037 FOT 53292 161/1820 N/A Unable to Load ATC
01-032 FOT 53358 166/0401 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-032 FOT see list June 01 14170 Generic Late Acquisition
01-033 FOT 53222 157/0220 N/A /CDEL function not working
01-034 FOT 53225 157/0754 N/A Unable to dump ATC
01-037 FOT 53292 161/1820 N/A Unable to load ATC
SEVNTFW01/24